OpenStudy (anonymous):

The lesson in my intermediate algebra math class is radical expressions. The problem I am having trouble on is (4y+1)^1/3 = -1 So far I understand that the fraction exponent is actually a cube root and for some reason, I keep getting y = -2/4

OpenStudy (anonymous):

if the cubed root is -1, then the whole thing is -1, i.e \[4y+1=-1\]

OpenStudy (anonymous):

if you keep getting \(-\frac{2}{4}\) that is good, but you might want to write it as \(-\frac{1}{2}\)

OpenStudy (anonymous):

Oh, okay. The thing is my teacher doesn't like fractions so I figured the answer shouldn't be a fraction. Thank you.